kekeman:

Factor x2 + 64. Check your work.

kekeman:

Factor x^2 + 64. Check your work.

Tranquility:

I'm sure you're aware of the differences of two squares: \( a^2 - b^2 = (a+b)(a-b)\) However, for your question you need to use \( a^2 + b^2 = (a+bi)(a-bi)\) i is the imaginary number \(\sqrt{-1}\)

kekeman:

Ok so (x+8i)(x-8i)

Tranquility:

That's correct!

kekeman:

Why is it not letting me close the post

UGKRza:

it did the same to me lol

kekeman:

Oof

Tranquility:

Seems like there's a bug with the close button. You can ask your question in another subject and then I can move it to mathematics for you

kekeman:

Ok